OpenStudy (bossimbacon):

which ordered pair is a solution to the system of equations? x+y=3 2x+2y=6

pooja195 (pooja195):

Hi there! Any ideas on where to start?

OpenStudy (bossimbacon):

umm i dont know exactly what to do

pooja195 (pooja195):

Well we need to find the value of x and y so... x+y=3 Can you attempt to isolate x? Tell me what equation you get. ^_^

OpenStudy (bossimbacon):

subtract y from both sides to get x=3y

OpenStudy (bossimbacon):

x= 3-y i mean sorry

pooja195 (pooja195):

Good! :) Now we have 2 equations x=3-y <--equation 1 2x+2y=6 <---equation 2 Take e1 since we know the value of x and substitute it into equation2 \[\huge~\rm~\bf~ 2(\color{purple}{3-y})+2y=6\]

OpenStudy (bossimbacon):

do i substitue it in for x or y?

pooja195 (pooja195):

I plugged it in for you, you need to solve for y

OpenStudy (bossimbacon):

ok so do i do the distributive prperty first

pooja195 (pooja195):

Yes

OpenStudy (bossimbacon):

6-2y+2y=6

OpenStudy (bossimbacon):

and then 2y + 2y = 4y and then 6-4y= 6

pooja195 (pooja195):

Wait -2+2=?

OpenStudy (bossimbacon):

ooo my bad, so is there just on y or no y's at all

OpenStudy (bossimbacon):

one y*

OpenStudy (bossimbacon):

would it be 6-y = 6

pooja195 (pooja195):

It would cancel out (2)(−y)+(2)(3)+2y=6(Distribute) −2y+6+2y=6 -2y+2y=0 6=6

pooja195 (pooja195):

We cannot solve from here. So our answer would be....?

OpenStudy (bossimbacon):

infinite?

pooja195 (pooja195):

Perfect! :D

OpenStudy (bossimbacon):

omg thank you so much! can tag you in some more questions later?

pooja195 (pooja195):

Of course you may ^_^ yw ^.^